OpenStudy (anonymous):

solve by factoring 2a^2 - 3a = -5

OpenStudy (anonymous):

2a^2-3a=-5 2a^2-3a+5=0 a^2-3a+10=0 (a+2)(a-5)=0 a+2=0 a=-2 a-5=0 a=5 5 and -2

OpenStudy (anonymous):

(a+2)(a-5)=0 => a=5,-2

jhonyy9 (jhonyy9):

2a2 -3a +5 not is equal (a+2)(a-5) so i think that this not is right,correct sure !!!

jhonyy9 (jhonyy9):

2a2 -3a+5=0 a1,2=(3+/- sqrt(9-40))/4 =(3+/- isqrt31)/4

jhonyy9 (jhonyy9):

,,BellaRue" before you accept one answer like right,correct you can check it to be sure that is the answer what you like wane getting on this exercise